### v2

@@ 5.6 Visualization
In this experiment, we perform a visualization to empirically compare the interpretability of selected features. The best results of three tensor-based methods and one non-tensor-based method in terms of NMI on two image datasets: JAFFE and BreastMNIST, are visualized in Figure 6, where we highlight the selected features to see the attention region of the UFS methods. We have the following observations:
@@
    STPCA-MP pays more attention to discriminative features. The exploitation of data orientation enables STPCA-MP to mine essential information with different spatial distributions. On JAFFE, STPCA-MP selects more horizontal features that cover the regions of eyes, eyebrows, mouth, and hair, etc. beneficial for distinguishing facial expressions, while CPUFS and MSPCA are misguided by background pixels. NLGMS, though it selects reasonable features as well, has a more scattered region and fails to select correlated features. On BreastMNIST, the features selected STPCA-MP tend to be grouped at an abnormal shadowed region composed of both horizontally and vertically distributed features, while the results of CPUFS, MSPCA and NLGMS are diverse. We can also observe that compared to CPUFS, MSPCA selects more rows (or columns) containing discriminative features and is less affected by the background, which we believe is attributed to the direct utilization of sparse factor matrices (CPUFS doesn’t apply a sparse tensor decomposition).
@@
    The score map of STPCA-MP exhibits significantly clearer and more distinct patterns than comparative methods. Thanks to the slice-by-slice operation, STPCA-MP describes feature correlations in a more delicate manner. On JAFFE, the score map of STPCA-MP approximates the shape of a human face. On BreastMNIST, the score map of STPCA-MP demonstrates a smooth score gradient transition radiating from the center of the shadowed region. In contrast, the score maps of CPUFS and MSPCA are less interpretable, due to the vague description of mode relations rooted in CPD and Tucker decomposition. As for NLGMS, its score map adds significant weights to the redundant features in the background. This phenomenon is possibly evoked by the sensitivity to noise features of the constructed graph.
Figure 6: Visualization on JAFFE and BreastMNIST. We show the score maps and highlight the positions of the top-ranked features in the samples. We set the number of selected features to $300$ on JAFFE, and $50$ on BreastMNIST.
@@ 5.7 Training time comparison
In this experiment, we compare the computational cost of comparative methods. The training time corresponding to the best clustering results is given in Table 4, while the average training time in terms of regularization parameter combinations is illustrated in Figure 7. We have the following observations:
@@
    Tensor-based methods are generally more efficient on high-dimensional data. By leveraging tensor-based techniques, the computational complexity is reduced from quadratic or cubic in the full feature space to quadratic or cubic within a single mode dimension. Therefore CPUFS, CPUFSnn, MSPCA, and our proposed method averagely cost less training time on JAFFE, UMCS, and ASCH compared to other methods.
@@
    STPCA-MP is efficient on both high-dimensional and large-scale data. CPUFS and CPUFSnn usually take hundreds of iterations to converge because of the non-convexity. Moreover, their computational complexity scales quadratically with #$\#(Sample)$, whereas STPCA-MP exhibits linear complexity relative to #$\#(Sample)$. Therefore, STPCA-MP takes considerably less training time than CPUFS and CPUFSnn on all datasets, especially on large-scale datasets. Although MSPCA is on average more efficient than STPCA-MP due to a smaller weight parameter size requiring computation in Tucker decomposition, it gains no advantage over the proposed method in both clustering tasks and the POC metric. Moreover, the training time associated with optimal performance shows only a marginal discrepancy between MSPCA and STPCA-MP (See Table 4), which is in an acceptable range. On the whole, the proposed method is both effective and efficient.
@@ 5.8 Convergence analysis
In this experiment, we study the convergence of the proposed method. Figure 8 illustrates the converge curves of objective function value on four real-world datasets. Utilizing the HPSD projector, the proposed two-step algorithm converges fast into the solution space within dozens of iterations.
@@
Figure 8: Converge curves of STPCA-MP on JAFFE and BreastMNIST. All regularization parameters are set to 1.
@@ 5.9 Parameter sensitivity analysis
In this section, we analyze the impact of the parameters $\lambda$ and $\eta$ in STPCA-MP on the performance of feature selection. We vary the parameters in the range of $\{10^{-2},10^{-1},1,10^{1},10^{2}\}$, set the selected features to 100, and record the ACC and NMI of STPCA-MP with all possible parameter combinations. The results on JAFFE and USPS is illustrated in Figure 9. We can see that
@@
    On JAFFE, the performance of STPCA-MP is relatively stable when $\lambda$ and $\eta$ vary in a certain range, and degenerates when the parameters become too large. This tendency is consistent between ACC and NMI.
@@
    On USPS, STPCA-MP is generally not sensitive to the choice of parameters, achieving relatively stable performance with respect to both ACC and NMI.
@@
5.10 Discussion on different choices on data orientation and transform domain.
In this experiment, we discuss the influence of data tensor orientation and transform domain. We conduct STPCA-MP with different data tensor rotations and transform matrices $\mathbf{M}$ on several datasets. The detailed ACC curves on JAFFE and USPS are illustrated in Figure 10. We further provide the best ACC and the mean POC in Table 5. The visualized result with $\mathbf{M}=\mathbf{I}$ on PIE and JAFFE is given in Figure 11. We have the following observations:
@@
    Different data orientations can significantly affect STPCA-MP’s clustering performance. We can see significant gaps among certain ACC curves in Figure 10. Like looking at an identical object from different views, decomposing a data tensor in different directions naturally leads to discrepant information. The visualized result in Figure 11 implies that STPCA-MP can capture features with different spatial distributions given different data orientations. A proper choice of data orientation is beneficial for mining feature correlations. The spatial distribution and latent correlation structure of discriminative features are heterogeneous among data tensors. On UCIDSA and UMCS, the overall performance of the Dir1s is better than the Dir2s, since the feature relations in multi-channel time-series data exist along 1-mode. In conclusion, performing slice-by-slice operations along a certain direction benefits UFS at multiple levels on different datasets.
@@
    A suitable transform domain can benefit the performance. Since the transform extracts the information from one of the two feature modes, while a slice-by-slice operation is conducted along the other feature mode, a proper transform domain can integrate more feature correlations as prior knowledge for the following optimization and UFS. In Table 5, Dir1-EigVector sometimes achieves better results than Dir-I, especially on time-series data. However, there is no absolute conclusion that certain settings of transform domain necessarily outperform others. As we can observe in both Figure 10 and Table 5, there is no setting under which the metrics consistently outperform all other configurations. Identical settings of orientation and transform domain can lead to substantial performance discrepancies across different datasets. In general, the choice of orientation and transform domains requires detailed configurations to achieve the best performance.
@@ 6 Conclusion
In this paper, we propose a novel Sparse Tensor PCA method named STPCA-MP by combining the encoder-decoder structure of SPCA and the orientation-dependent framework of T-SVDM. We deduce a convex version of STPCA-MP that utilizes the sparse reconstruction tensor to perform UFS. An efficient slice-by-slice algorithm to conduct STPCA-MP is proposed to guarantee fast convergence. The experimental results on real-world data show that our method outperforms state-of-the-art methods on average and can handle different data organization scenarios at a smaller training time cost. In particular, STPCA-MP has the advantage of selecting discriminative features with the greatest BCV on slice-wise data. In addition, we also test the parameter sensitivity, examine the efficient convergence, and discuss the impacts of choosing different data orientations and transform domains. The experimental results show that with a suitable combination of data orientation and transform domain, our method can capture abundant discriminative information and achieve remarkable performance.
There are two valuable research points concerning STPCA-MP in the future. First, the best choice of transform domain is possibly related to the background knowledge of data. More experiments should be carried out for further applications. Second, the extension to higher-order scenarios of STPCA-MP still lacks an effective and efficient optimization algorithm. We will be working on addressing the above two problems in the future.
Appendix A Appendix
@@
Thus, we can rewrite the objection function as:

### v3

@@ 5.6 Visualization
In this experiment, we conduct feature visualization to empirically compare the interpretability of selected features. Figure 7 displays the best NMI results from three tensor-based methods (STPCA-MP, CPUFS, MSPCA) and one non-tensor method (NLGMS) on three image datasets: JAFFE, BreastMNIST, and PIE, with selected features highlighted to reveal method attention patterns. We have the following observations:
@@
    STPCA-MP demonstrates superior discriminative feature selection. By leveraging data orientation, STPCA-MP effectively captures essential spatial patterns: (i) On PIE and JAFFE, it consistently selects horizontal features corresponding to facial components (eyes, eyebrows, mouth, hair) crucial for facial and expression recognition (Figure 7(b)(k)), while CPUFS, NLGMS and MSPCA misallocate attention to background pixels (Figure 7(c)(l)(d)(m)); (ii) On BreastMNIST, it concentrates on relevant shadowed regions with both horizontal and vertical features (Figure 7(f)), unlike the scattered selections of CPUFS, MSPCA, and NLGMS (Figure 7(g)(h)(I)). Notably, MSPCA’s sparse tensor decomposition provides moderate improvement over CPUFS by reducing background interference (Figure 7(d)(h)(m)).
@@
    STPCA-MP produces more interpretable score maps. The slice-by-slice operation yields precise feature correlation modeling on: (i) Facial datasets, where STPCA-MP exhibits anatomically plausible attention maps resembling face contours (Figure 7(b)(k)); (ii) BreastMNIST, where STPCA-MP shows smooth score gradients radiating from pathological regions (Figure 7(f)). In contrast, CPUFS and MSPCA generate less coherent patterns due to inherent limitations of CPD/Tucker decompositions in modeling mode relations (Figure 7(c)(d)(g)(h)(l)(m)), while NLGMS overweight background noise (Figure 7(e)(i)), likely due to graph construction sensitivity.
Figure 7: Visualization on PIE, JAFFE, and BreastMNIST. We show the score maps and highlight the positions of the top-ranked features in the samples. We set the number of selected features to $300$ on JAFFE, $50$ on BreastMNIST, and $100$ on PIE.
@@ 5.7 Training time comparison
In this experiment, we analyze the computational efficiency of comparative methods. The training time corresponding to optimal clustering performance is reported in Table 7, with average training times across parameter combinations shown in Figure 8. We have the following observations:
@@
    Tensor-based methods demonstrate superior efficiency on high-dimensional data. By exploiting tensor operations, these methods reduce computational complexity from quadratic/cubic in full feature space to quadratic/cubic within single-mode dimensions. This explains why CPUFS, CPUFSnn, MSPCA, and our method consistently achieve lower training times on JAFFE, UMCS, and ASCH datasets compared to non-tensor alternatives (Figure 8(a)(b)(c)).
@@
    STPCA-MP maintains efficiency across both high-dimensional and large-scale datasets.While CPUFS and CPUFSnn require hundreds of iterations (due to non-convexity) with quadratic sample complexity $\mathcal{O}(n^{2})$, STPCA-MP achieves linear complexity $\mathcal{O}(n)$. This efficiency advantage becomes particularly pronounced on large-scale datasets like USPS (Figure 8(d)). Although MSPCA’s Tucker decomposition with smaller weight parameters yields marginally faster average training times, it fails to outperform STPCA-MP in either clustering metrics or POC (Table 7). Notably, the runtime difference between MSPCA and our method at optimal performance is negligible (Table 7), confirming STPCA-MP’s balanced effectiveness-efficiency trade-off.
@@ 5.8 Convergence analysis
In this experiment, we investigate the convergence properties of the proposed method through a representative subproblem analysis due to space constraints. Figure 9 demonstrates the convergence behavior by plotting both the objective function values and runtime per iteration across four real-world datasets. The two-step algorithm achieves rapid convergence to the solution space within dozens of iterations, facilitated by the HPSD projector that ensures stable optimization. The slice-by-slice operation significantly reduces computational complexity, enabling efficient execution of each iteration while maintaining solution quality.
@@
Figure 9: Convergence curves and runtime per iteration of STPCA-MP on four real-world datasets. All regularization parameters are set to 10.
@@ 5.9 Parameter sensitivity analysis
In this section, we evaluate the impact of regularization parameters $\lambda$ and $\eta$ on the feature selection performance of STPCA-MP. By varying both parameters within $\{10^{-2},10^{-1},1,10^{1},10^{2}\}$ while fixing the selected features at 100, we record the ACC and NMI values across all parameter combinations. Figure 10 presents the results on JAFFE and USPS datasets, revealing two key findings:
@@
    For JAFFE, STPCA-MP maintains stable performance when $\lambda$ and $\eta$ remain within a moderate range, with noticeable degradation occurring only at excessively large parameter values - a trend consistently observed in both ACC and NMI metrics.
@@
    For USPS, the method demonstrates robust performance across the tested parameter range, showing minimal sensitivity to $\lambda$ and $\eta$ variations in terms of both ACC and NMI measurements.
@@
5.10 Impact of Data Orientation and Transform Domain
In this experiment, investigate how data tensor orientation and transform matrix $\mathbf{M}$ affect STPCA-MP performance. Experiments evaluate different tensor rotations and transform matrices across multiple datasets, with detailed ACC curves for JAFFE and USPS shown in Figure 11. Table 8 reports the optimal ACC and mean POC values, while Figure 12 visualizes results using $\mathbf{M}=\mathbf{I}$ on PIE and JAFFE. We have the following observations:
@@ 5.10 Impact of Data Orientation and Transform Domain
    Data orientation critically influences clustering performance. Significant ACC variations in Figure 11(a) and (b) demonstrate that, similar to viewing objects from different perspectives, tensor decomposition along distinct directions captures varying information. Figure 12 confirms STPCA-MP’s ability to identify orientation-dependent spatial feature distributions. Optimal orientation selection enhances feature correlation mining, particularly for multi-channel time-series data like UCIDSA and UMCS, where 1-mode operations outperform due to inherent feature relationships along this dimension.
@@
    Transform domain selection impacts results. A well-selected transform matrix extracts information from one feature mode while slice-by-slice operations process the other, enabling prior knowledge incorporation for optimization. Table 8 shows Dir1-EigVector occasionally surpassing Dir-I, especially for time-series data. However, no transform configuration universally dominates - both Figure 11(a)(b)(c)(d) and Table 8 demonstrate performance varies substantially across datasets. Therefore, optimal orientation and transform choices require dataset-specific tuning.
@@ 6 Conclusion
In this paper, we propose a novel Sparse Tensor PCA method named STPCA-MP by combining the encoder-decoder structure of SPCA and the orientation-dependent framework of T-SVDM. We deduce a convex version of STPCA-MP that utilizes the sparse reconstruction tensor to perform UFS. An efficient slice-by-slice algorithm to conduct STPCA-MP is proposed to guarantee fast convergence. The experimental results on real-world data show that our method outperforms state-of-the-art methods on average in terms of clustering metrics (ACC, NMI, and SS) in different data organization scenarios at a smaller training time cost. In particular, STPCA-MP has the advantage of selecting discriminative features with the greatest BCV on slice-wise data. In addition, we also test the parameter sensitivity, examine the efficient convergence, and discuss the impacts of choosing different data orientations and transform domains. The experimental results show that when the transform axes align with the feature distribution pattern, our method can capture abundant discriminative information and achieve remarkable performance.
There are two valuable research points concerning STPCA-MP for the future. First, the best choice of transform domain is possibly related to the data background knowledge. More experiments should be carried out for further applications. Second, the extension to higher-order scenarios of STPCA-MP still lacks an effective and efficient optimization algorithm. We will be working on addressing the above two problems in the future.
Appendix A Proof of Theorem 1
@@
Thus, we can rewrite the objective function as:
@@
Appendix B Reproducibility Statement
To facilitate the reproducibility of our research, we provide the following resources:
Code Implementation
  • Code Repository: Available at https://github.com/zjj20212035/STPCA.git under the main branch (git checkout v1.1).
  • Dependencies:
    • MATLAB R2023b
    • Tensor Toolbox v3.5 (included in utils/)
  • Parallel Computing: All experiments utilize MATLAB’s Parallel Computing Toolbox.
Datasets
  • All datasets are provided in data/ folder as .mat files, each is unified to contain the following common attributes:
    • data: $d_{1}\times d_{2}\times n$ tensor ($n$=samples, $d_{i}$=$i$-mode dimensionality)
    • label: $n\times 1$ class vector
    • tensor_type: tube-wise or slice-wise
    • tensor_size: dimensionality of the feature space
  • No manual data preprocessing needed - all preprocessing is handled automatically by:
    • config_Table7.m
    • config_Figure7.m
    where each dataset shares the same normalization that is carried out by dividing each element by the maximum absolute value, scaling all values to the interval $[-1,1]$.
Experiment Reproduction
We have streamlined the code to allow one-click reproduction of the paper’s key results. Open the project in MATLAB and
  • To replicate Table 7:
        cd code
        run(’exp_Table7_clustering_POC.m’)
  • To replicate Figure 7:
        cd code
        run(’exp_Figure7_visualization.m’)
Hyperparameters are specified in config_Table7.m and config_Figure7.m, where:
  • Regularization parameters $\lambda$ and $\eta$ are tuned by grid search in the range of $\{10^{-2},10^{-1},1,10,10^{2}\}$,
  • Orientation is tuned from $D^{o}=\{\{1,3,2\},\{2,3,1\}\}$ and the best result is shown,
  • The stabilization constants are fixed at $\epsilon_{1}=10^{-3}$ and $\epsilon_{2}=10^{-2}$ throughout all experiments.